OpenStudy (anonymous):

Simplify i^42 simplify i^72

OpenStudy (anonymous):

I think the answer for both is -1, it is just like saying i^2, which is equal to = -1

OpenStudy (anonymous):

sorry I got it now, i^42 = i*i*i*i*... 42 times... correct? now we know that i*i= -1 so how many i*i do we have from the exponent 42? 21 pairs, which means 21 (-1)... correct? so -1^21= -1 hence, i^42 = -1 Now for 72 we have 36 pairs of i*i, therefore 36 of (-1), which means (-1)^36 = 1, and hence i^72 = +1 do you get it? let me know if you have any questions! =)